X 5.3.25 Question Suppose Emerson wins 37% of all bingo games (a) What is the probability that Emerson wins two bingo games in a row? (b) What is the probability that Emerson wins three bingo games in a row? (c) When events are independent, their complements are independent as well. Use this result to determine the probability that Emerson wins three bingo games in a row, but does not win four in a row. (a) The probability that Emerson wins two bingo games in a row is

Suppose Emerson wins 37% of all bingo games 

(a) What is the probability that Emerson wins two bingo games in a row? 

(b) What is the probability that Emerson wins three bingo games in a row?

(c) When events are independent, their complements are independent as well. Use this result to determine the probability that Emerson wins three bingo games in a row, but does not win four in a row. 

(a) The probability that Emerson wins two bingo games in a row is
Answer #1

a)P(wins two bingo in a row)=0.37*0.37=0.1369

b)P(wins three bingo in a row)=0.37*0.37*0.37=0.050653

c) P(not wins four in a row but not wins four in a row =0.37*0.37*0.37*(1-0.37)=0.031911